OpenStudy (anonymous):

Anita has a collection of sports trading cards. Her collection consists of three basketball cards, six baseball cards, seven hockey cards, and eight football cards. Anita picked one card from her collection. What is the probability that Anita picked a football card?

OpenStudy (anonymous):

8/total possibilities =8/(3+6+7+8) =8/24 =1/3